Given Z=f(x-2y)+g(3x+y) such that Zxx + AZxy+BZyy=0 Find A and B
1
Expert's answer
2020-03-03T09:19:07-0500

Let's denote the expression "x-2y=\zeta (x,y)" and "3x+y=\xi(x,y)". Then rewrite task as

(1) "Z(x,y)=f(\zeta(x,y))+g(\xi(x,y))"

The first derivatives by the rule of differentiation of nested functions will be

(2) "Z'_x=f'_{\zeta}\cdot \zeta'_x+g'_{\xi}\cdot \xi'_x=f'_{\zeta}+3g'_{\xi}"

(3) "Z'_y=f'_{\zeta}\cdot \zeta'_y+g'_{\xi}\cdot \xi'_y=-2f'_{\zeta}+g'_{\xi}"

The second derivatives are

(4) "Z''_{xx}=f''_{\zeta\zeta}\cdot \zeta'_x+3g''_{\xi\xi}\cdot \xi'_x=f''_{\zeta\zeta}+9g''_{\xi\xi}"

(5) "Z''_{xy}=f''_{\zeta\zeta}\cdot \zeta'_y+3g''_{\xi\xi}\cdot \xi'_y=-2f''_{\zeta\zeta}+3g''_{\xi\xi}"

(6) "Z''_{yy}=-2f''_{\zeta\zeta}\cdot \zeta'_y+g''_{\xi\xi}\cdot \xi'_y=4f''_{\zeta\zeta}+g''_{\xi\xi}"

Let's substitute these expressions into the equation of the problem and give similar terms.

(7) "Z''_{xx}+AZ''_{xy}+BZ''_{yy}=0\\f''_{\zeta\zeta}+9g''_{\xi\xi}+A\cdot (-2f''_{\zeta\zeta}+3g''_{\xi\xi})+B\cdot(4f''_{\zeta\zeta}+g''_{\xi\xi})=0"

(8) "f''_{\zeta\zeta}\cdot (1-2A+4B)+g''_{\xi\xi}\cdot(9+3A+B)=0"

Equation (8) can be performed for any "f" and "g" only if the expressions in parentheses are both 0. We get a system of two algebraic equations with two unknowns.

(9) "\begin{cases}
 1-2A+4B=0 \\
 9+3A+B=0
\end{cases}"

In standard form they are

(10) "\begin{cases}
 2A-4B=1 \\
 3A+B=-9
\end{cases}"

The solution to this system is easy to find "A=-\frac{35}{14}; B=-\frac{21}{14}" or "A=-\frac{5}{2}; B=-\frac{3}{2}"

Answer: "A=-\frac{5}{2}; B=-\frac{3}{2}"